# Emotion-Coherent Reasoning for Multimodal LLMs via Emotional Rationale Verifier

### v2

@@ 3.3 Explanation Reward
Final reward.
In addition to the explanation reward $R_{E}$, we incorporate two auxiliary rewards format reward and answer reward to compute the final total reward $R$. The format reward $R_{\text{format}}$, following prior work such as RLVR, assigns 1 if the model output adheres to the required format (i.e., includes both <think> and <answer> tags), and 0 otherwise. The answer reward $R_{\text{answer}}$ assigns 1 if the predicted emotion within the <answer> section matches the GT emotion label $e_{\text{gt}}$.
The final reward used in training is computed as:
@@
By incorporating the explanation reward $R_{E}$ obtained from ERV as an additional learning signal during GRPO training, we aim to enhance the consistency between the model’s reasoning process and the GT emotion.
@@
Human-Omni (Zhao et al. 2025b)     - 20.18 13.52     - 22.64 19.44
@@
Emotion-LLaMA (7B) (Cheng et al. 2024a)     - - - - -     41.15 32.56 57.69 48.71 38.54
@@ 4.1 Dataset
EMER (Lian et al. 2023) is a human-annotated emotional video dataset with 332 samples labeled with one of five emotions (angry, sad, surprise, worried, happy) and corresponding descriptions. We use EMER to train the HumanOmni model to generate explanations in a reasoning format.
MERR (Cheng et al. 2024a) provides 4,487 fine-grained pseudo-labeled video-caption pairs, generated via facial and audio analysis with LLM-guided annotation. We use this data to supervise our ERV, leveraging its diverse and component-level emotion cues.
DFEW (Jiang et al. 2020) is a large-scale facial expression video dataset with 9,362 training and 2,342 test samples, labeled with 7 basic emotions: angry, disgust, surprise, happy, sad, neutral, and fear (official 5th split). We use DFEW in two ways: (1) generating textual explanations from training videos and assigning pseudo-labels via GPT-4.1 to augment ERV training, and (2) applying our reward-based RL scheme using EMER-pretrained models to produce <think> and <answer> outputs.
MAFW (Liu et al. 2022) is a multimodal emotion dataset with 7 basic and 4 compound categories (contempt, helplessness, anxiety, disappointment), comprising 7,341 training and 1,831 test samples (set 5 split). We apply the same pipeline as DFEW, pseudo-labeled explanation augmentation for ERV and RL training using our proposed rewards.
@@ 4.2 Evaluation Metrics
Emotion Recognition Accuracy Following prior works (Zhao, Wei, and Bo 2025), we employ Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) to measure the model’s emotion classification performance.
Evaluating Emotional Coherence We evaluate the emotional coherence of generated explanations using three metrics: Explanation Emotion Accuracy (EEA), Explanation–Prediction Consistency (EPC), and Faithful Consistency Rate (FCR). These metrics capture different aspects of alignment among the explanation, predicted emotion, and ground-truth label. GPT-4.1-mini is used to recognize emotion in explanation. Figure 3 shows the prompt used in our evaluation. The detailed formulations of these metrics are provided in Appendix C.
@@ 4.3 Implementation Details
We employ HumanOmni 0.5B model (Zhao et al. 2025b) as a backbone model. Our ERV is based on the RoBERTA-100M trained on the GoEmotions Dataset. Our training pipeline consists of three components: SFT, ERV module, and GRPO training. The main model is first fine-tuned on the EMER dataset for 5 epochs with a learning rate of $2\mathrm{e}^{-5}$, a cosine scheduler, a warmup ratio of 0.03, and a batch size of 32. The ERV module is trained separately for 10 epochs with a batch size of 64 and a learning rate of $3\mathrm{e}^{-5}$. Those are trained on 8 NVIDIA RTX 3090 GPUs. For GRPO training, we use a batch size of 16 with a gradient accumulation step of 2, generating $G=16$ samples per input, while applying a KL divergence penalty with coefficient $\beta=0.04$ and fixing the sampling temperature at 0.3 during inference. For the 7B model, we trained with $G=4$ instead. This stage is conducted on 8 NVIDIA A100 GPUs. Appendix E shows more training details.
@@ 5.1 Main Results
To verify the effectiveness of our proposed Explanation reward in the quality of generated explanations, we evaluated on three proposed metrics: EEA, EPC, and FCR, with standard recognition accuracy WAR and UAR. Our model is compared with three models trained in different ways using the same multimodal emotion description dataset and HumanOmni architecture. (1) EMER-SFT: A supervised fine-tuned model trained on the EMER dataset for emotion reasoning, available in both 0.5B and 7B LLM versions. (2) R1-Omni: A model trained with the GRPO using answer and format rewards. (3) Baseline: Since R1-Omni lacks training details and relies on an in-house dataset, we re-trained EMER-SFT with GRPO on the answer and format rewards using the same training dataset and hyperparameter settings as our method for a fair comparison.
@@ 5.2 Ablation Study
Table 2: Ablation study on the Emotional Rationale Verifier (ERV) and Explanation Reward on MAFW
@@
To verify our proposed training strategy on the ERV module, we removed the data augmentation stage utilizing GPT-4.1. (1) Baseline, without the ERV module, can’t reward on the generated explanation. It only focuses on the format and final answer. (2) Without a data augmentation strategy, training the ERV solely on the EMER and MERR-fine datasets boosts EEA from 32.41% to 46.88% and FCR from 27.27% to 42.62%, while keeping WAR and UAR unchanged. (3) Adding the augmented data to ERV further raises EEA + 1.26 pp, FCR + 1.31 pp, EPC + 2.90 pp, and overall recognition accuracy. These results suggest that the ERV’s judgment capability is constrained by the size of its training data. When additional multimodal-emotion descriptions are unavailable, text-level data augmentation noticeably enhances the model’s multimodal reasoning ability.
@@
To validate the effectiveness of our proposed Explanation Reward, we compare it against a variant trained without neutral filtering in the reward computation. That is, the reward does not take into account the proportion of non-emotional sentences (as described in Equation 4 for the neutral case). Without neutral filtering, the overall performance drops by 2.65 pp on EEA, 1.44 pp on FCR, 3.88 pp on EPC, 0.33 pp on WAR, and 0.45 pp on UAR. Furthermore, when sentence-level explanation verification is removed, we also observed performance degradation, with EEA and FCR dropping by 1.39 pp and 0.40 pp, respectively, and recognition metrics (WAR and UAR) also decreasing by 0.71 pp and 0.46 pp. As a result, when providing emotional guidance for explanation, it is effective to provide rewards according to emotional sentence proportion at the sentence level.
5.3 Human Evaluation about Explanation Quality
Through the main results and ablation studies, we confirmed that explanations from our model better align with both ground-truth and predicted emotions. However, their quality in terms of how well they explain the emotions in the videos had not been directly evaluated. To address this, we conducted a human study comparing our model’s explanations to those from R1-Omni. Specifically, we randomly selected 14 samples from each model (total 28), where the generated emotion matched the ground-truth. Twenty human raters evaluated each explanation on a 1–5 scale based on how well it explained the corresponding video’s emotion. For each sample, we recorded which model received the higher score and computed the average score across all samples. As shown in Table 3, our model achieved a 53.6% win rate and a higher average score (3.52 vs. 3.42), indicating that it produces more informative and emotionally grounded explanations than R1-Omni. (Survey details are provided in Appendix A.)
@@ 5.3 Human Evaluation about Explanation Quality
Table 4: Comparison of Baseline and Ours across different closed-source LLMs in terms of EEA, FCR, and EPC.
@@ 5.4 Robustness of the Proposed Reasoning Evaluation Metrics
Since the evaluation of EEA, FCR, and EPC was conducted using a single closed-source LLM (GPT-4.1-mini), we further validated the robustness of the proposed metrics by performing additional evaluations with a different model family, Gemini-2.5-Flash (Comanici et al. 2025). As shown in Table 4, the evaluation on the MAFW for both the Baseline and our model exhibits the same tendency of improved reasoning quality. This demonstrates that our metrics are independent of the specific LLM used for evaluation and consistently capture the alignment between reasoning and the expressed emotion across different model families. The small absolute differences (0.17–2.62 percentage points, approximately 0.5–3.8% relative deviation) further confirm the robustness and stability of the evaluation across measurement settings. Moreover, in the human survey, participants were asked to judge the emotion conveyed by the model-generated reasoning, and their judgments showed a 92.9% agreement with those of GPT-4.1-mini (see Appendix A).
@@ 6 Conclusion
We proposed training an MLLM for the MER task utilizing our proposed ERV and Explanation Reward, aiming to enhance the coherence between emotion recognition and explanation. Compared to prior approaches, where coherence between predictions and explanations among correctly predicted samples remained around half (55.13% on MAFW and 42.92% on DFEW), our method significantly improves this alignment, achieving 75.79% and 79.87%. Furthermore, human evaluations confirmed that our approach produces more emotionally and contextually grounded emotion explanations for video data. As a result, our work enables emotionally coherent and trustworthy explanations alongside accurate emotion recognition, representing a significant advancement toward emotionally intelligent human–computer interaction.
@@ A.1 Survey Questions
We provided three questions for each sample, as show in Figure 5 and 4.

### v3

@@ 3.3 Explanation Reward
Total reward.
In addition to the $R_{E}$, we incorporate two auxiliary rewards format reward $R_{\text{format}}$ and answer reward $R_{\text{answer}}$ to compute the total reward $R$. $R_{\text{format}}$, following prior work (guo2025deepseek), assigns 1 if the model output adheres to the required format (i.e., includes both <think> and <answer> tags), and 0 otherwise. $R_{\text{answer}}$ assigns 1 if the predicted emotion within the <answer> matches $e_{\text{gt}}$, and 0 otherwise. The total reward used in training is computed as:
@@
Incorporating $R_{E}$ during GRPO training enhances consistency between the model’s reasoning and the target emotion.
@@
Human-Omni (zhao2025humanomni)     - 20.18 13.52     - 22.64 19.44
@@
Emotion-LLaMA (7B) (cheng2024MERR)     - - - - -     41.15 32.56 57.69 48.71 38.54
@@ 4.1 Dataset
EMER (lian2023EMER) is a human-annotated emotional video dataset with 332 samples labeled with one of five emotions (angry, sad, surprise, worried, happy) and corresponding descriptions. It is used to train the model to generate explanations in a reasoning format.
MERR (cheng2024MERR) provides 4,487 fine-grained pseudo-labeled video-caption pairs, generated via facial and audio analysis with LLM-guided annotation. We use this data to supervise our ERV, leveraging its diverse and component-level emotion cues.
DFEW (jiang2020dfew) is a large-scale facial expression video dataset with 9,362 training and 2,342 test samples, labeled with 7 basic emotions: angry, disgust, surprise, happy, sad, neutral, and fear (official 5th split). We use DFEW in two ways: (1) generating textual explanations from training videos and assigning pseudo-labels via GPT-4.1 to augment ERV training, and (2) applying our reward-based RL scheme.
MAFW (liu2022mafw) is a multimodal emotion dataset with 7 basic and 4 compound categories (contempt, helplessness, anxiety, disappointment), comprising 7,341 training and 1,831 test samples (set 5 split). We apply the same pipeline as DFEW.
@@ 4.2 Evaluation Metrics
Emotion Recognition Accuracy Following prior works (zhao2025r1), we employ Unweighted Average Recall (UAR) and Weighted Average Recall (WAR) to measure the model’s emotion classification performance.
Evaluating Emotional Coherence We evaluate the emotional coherence of generated explanations using three metrics: Explanation Emotion Accuracy (EEA), Explanation–Prediction Consistency (EPC), and Faithful Consistency Rate (FCR). These metrics capture different aspects of alignment among the explanation, predicted emotion, and ground-truth label. GPT-4.1-mini is used to recognize emotion in explanation. Figure 3 shows the prompt used in evaluation.
@@ 4.3 Implementation Details
We employ the HumanOmni 0.5B model (zhao2025humanomni) as our backbone and RoBERTA-100M (trained on GoEmotions) for the ERV. Our training pipeline consists of three components: SFT, ERV, and GRPO. The main model is first fine-tuned on the EMER dataset for 5 epochs with a $2\mathrm{e}^{-5}$ learning rate, a cosine scheduler, a 0.03 warmup ratio, and a batch size of 32. The ERV module is trained separately for 10 epochs with a batch size of 64 and a $3\mathrm{e}^{-5}$ learning rate. Those are trained on 8 NVIDIA RTX 3090 GPUs. For GRPO training, we use a batch size of 16 with 2 gradient accumulation steps, generating $G=16$ samples per input, and apply a KL divergence penalty of $0.04$. For inference, we use a temperature of 0.3. For the 7B model, we trained with $G=4$ instead. This stage is conducted on 8 NVIDIA A100 GPUs. Appendix E shows more training details.
@@ 5.1 Main Results
To verify the effectiveness of our proposed Explanation reward in the quality of generated explanations, we evaluated on three proposed metrics: EEA, EPC, and FCR, with standard recognition accuracy WAR and UAR. Our model is compared with three models trained in different ways using the same multimodal emotion description dataset and HumanOmni architecture. (1) EMER-SFT: A supervised fine-tuned model trained on the EMER dataset for emotion reasoning, available in both 0.5B and 7B. (2) R1-Omni: A model trained with the GRPO using answer and format rewards. (3) Baseline: Since R1-Omni lacks training details and relies on an in-house dataset, we re-trained EMER-SFT with GRPO on the answer and format rewards using the same training dataset and hyperparameter settings as our method for a fair comparison.
@@
Table 2: Ablation study of the ERV and $R_{E}$ on MAFW
5.2 Ablation Study
@@
To verify our proposed training strategy on the ERV module, we removed the data augmentation stage. (1) Baseline, without the ERV module, can’t reward on the generated explanation. It only focuses on the format and answer. (2) Without a data augmentation strategy, training the ERV solely on the EMER and MERR-fine datasets boosts EEA from 32.41% to 46.88% and FCR from 27.27% to 42.62%, while keeping WAR and UAR unchanged. (3) Adding the augmented data to ERV further raises EEA + 1.26 pp, FCR + 1.31 pp, EPC + 2.90 pp, and overall recognition accuracy. These results suggest that the ERV’s judgment capability is constrained by the size of its training data. When additional multimodal-emotion descriptions are unavailable, text-level data augmentation noticeably enhances the model’s multimodal reasoning ability.
@@
To validate the effectiveness of our proposed Explanation Reward, we compare it against a variant trained without our two key components: neutral filtering and sentence-level computation. In this variant, the reward calculation does not filter out non-emotional sentences (as described in Equation 4 for the neutral case), and the ERV evaluates the emotional coherence of the entire explanation holistically. We observed a performance drop when each of these components was removed individually. Specifically neutral filtering, which excludes emotionally irrelevant sentences, was particularly effective in guiding the model to generate emotion-focused explanations.
@@
Table 4: Performance of Baseline and Ours (EEA, FCR, EPC) using different closed-source LLMs.
5.3 Human Evaluation about Explanation Quality
Through the main results and ablation studies, we confirmed that explanations from our model better align with both target and predicted emotions. However, their quality in terms of how well they explain the emotions in the videos had not been directly evaluated. To address this, we conducted a human study comparing our model’s explanations to those from R1-Omni. Specifically, we randomly selected 14 samples from each model (total 28), where the generated emotion matched the ground-truth. Twenty human raters evaluated each explanation on a 1–5 scale based on how well it explained the corresponding video’s emotion. For each sample, we recorded which model received the higher score and computed the average score across all samples. As shown in Table 3, our model achieved a 53.6% win rate and a higher average score (3.52 vs. 3.42), indicating that it produces more informative and emotionally grounded explanations than R1-Omni. (Survey details are provided in Appendix A.)
@@ 5.4 Robustness of the Proposed Reasoning Evaluation Metrics
Since the evaluation of EEA, FCR, and EPC was conducted using a single closed-source LLM, we further validated the robustness of the proposed metrics by repeating the evaluation with a different model family, Gemini-2.5-Flash (comanici2025gemini). As shown in Table 4, the evaluation on MAFW reveals a consistent trend, with both the Baseline and our model exhibiting improved reasoning quality. This demonstrates that our metrics are independent of the specific LLM evaluator and consistently capture the alignment between reasoning and emotion across model families. Moreover, a human survey confirmed a 92.9% agreement between participants’ judgments on the conveyed emotion and those of GPT-4.1-mini. (See Appendix A.)
@@ 6 Conclusion
We proposed training an MLLM for the MER task utilizing our proposed ERV and Explanation Reward, aiming to enhance the coherence between emotion recognition and explanation. Compared to prior approaches, where coherence between predictions and explanations among correctly predicted samples remained around half (55.13% on MAFW and 42.92% on DFEW), our method significantly improves this alignment, achieving 75.79% and 59.77%.Furthermore, human evaluations confirmed that our approach produces more emotionally and contextually grounded emotion explanations for video data. As a result, our work enables emotionally coherent and trustworthy explanations alongside accurate emotion recognition, representing a significant advancement toward emotionally intelligent human–computer interaction.
Acknowledgments
This work was partly supported by two funds: Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No. RS-2025-25442384, QuestioningAI: Development of Active Reasoning Cooperative Multimodal Agent Technology), and the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529).
@@ A.1 Survey Questions
We provided three questions for each sample, as show in Figure 4 and  5.